A spider spins a web with silk threads with a linear density of 1.75 µg/m .   A typical tension in the radial threads of such a web is 4 mN . Suppose a fly hits this web.What is the speed of disturbance traveling along the radial thread of the web? Express your answer to two significant figures and include the appropriate units.
Answer #1

Linear density, u = 1.75 x 10^-6 x 10^-3 = 1.75 x 10^-9 kg/m

Velocity, v = sqrt(T/u)

V = sqrt (0.004/1.75 x 10^-9) = 1.5 x 10^3 m/s

Comment in case any doubt please rate my answer....